# Federated Unsupervised Semantic Segmentation

### v1

@@ Abstract
This work explores the application of Federated Learning (FL) in Unsupervised Semantic image Segmentation (USS). Recent USS methods extract pixel-level features using frozen visual foundation models and refine them through self-supervised objectives that encourage semantic grouping. These features are then grouped to semantic clusters to produce segmentation masks. Extending these ideas to federated settings requires feature representation and cluster centroid alignment across distributed clients—an inherently difficult task under heterogeneous data distributions in the absence of supervision. To address this, we propose FUSS (Federated Unsupervised image Semantic Segmentation) which is, to our knowledge, the first framework to enable fully decentralized, label-free semantic segmentation training. FUSS introduces novel federation strategies that promote global consistency in feature and prototype space, jointly optimizing local segmentation heads and shared semantic centroids. Experiments on both benchmark and real-world datasets, including binary and multi-class segmentation tasks, show that FUSS consistently outperforms local-only client trainings as well as extensions of classical FL algorithms under varying client data distributions. To support reproducibility, full code will be released upon manuscript acceptance.
Federated Unsupervised Image Segmentation
@@ 1 Introduction
Semantic image segmentation is a fundamental computer vision task, serving as a core component of visual perception systems deployed in safety-critical applications like autonomous driving (Tang et al., 2022), industrial inspection (Psarras et al., 2024), and medical diagnostics (Chen et al., 2024). Unsupervised Semantic Segmentation (USS) has emerged as a family of methods that do not require the notoriously famous pixel-wise annotations. Recent USS approaches employ pretrained Convolutional Neural Networks (CNNs) (He et al., 2016) or Vision Foundation Models (VFMs) (Caron et al., 2021; Dosovitskiy et al., 2020) to extract pixel-wise feature embeddings. These representations are fed into learnable segmentation heads that are fine-tuned through self-supervised learning objectives such as contrastive losses (Guermazi et al., 2024), energy minimization (Hamilton et al., 2022), or eigenvector alignment (Kim et al., 2024). Unsupervised segmentation is ultimately achieved by simply clustering the learned representations.
@@
In safety-critical domains, strict privacy constraints on data storage and sharing introduce significant challenges in effectively training robust and generalizable DNN-based image segmentation models. To this end, Federated Learning (FL) (McMahan et al., 2017; Caldarola et al., 2022; Gong et al., 2021; Fantauzzo et al., 2022) has emerged as a promising paradigm that enables collaborative model training with decentralized data. Perhaps the most recognizable FL method is FedAvg McMahan et al. (2017), a method that simply aggregates the weights between federation clients. The simplicity and elegance of FedAvg render it applicable to almost every supervised or unsupervised training (Lubana et al., 2022; Liao et al., 2024) scenario, consisting FedAvg the standard benchmark in most FL applications.
In semantic image segmentation, FL has only been applied in the supervised training case (Miao et al., 2023). The special challenge of federated semantic image segmentation is that data heterogeneity across clients is coupled with large within-class variance due to fine-graining (i.e., pixel regions of semantically similar objects may differ significantly). Supervised training objectives (Miao et al., 2023) are known to alleviate such issues, since they can "push" the representations of similar objects to the same target values across all clients. Nevertheless, these questions remain open and unexplored for the Federated Unsupervised Semantic Segmentation problem, where such supervised signals simply cannot be obtained.
@@
To bridge this gap, we propose FUSS (Federated Unsupervised image Semantic Segmentation), a novel framework for collaborative unsupervised training of image segmentation DNN models in decentralized settings. FUSS addresses both data annotation scarcity and privacy constraints by enabling clients to jointly learn DNN segmentation heads and class prototypes using unsupervised objectives and communication-efficient parameter exchanges.
@@
  • We identify a fundamental gap in current federated image segmentation literature, namely, its reliance on annotated image data, which limits applicability in privacy-constrained data settings.
  • We introduce FUSS, a label-free federated learning framework that combines unsupervised representation learning with clustering-based semantic aggregation to enable efficient federated image segmentation training without human supervision.
@@
  • We develop a novel aggregation strategy, namely Federated Centroid Clustering (FedCC), tailored to federated unsupervised segmentation. FedCC supports joint optimization of client-specific segmentation heads and global alignment of semantic class prototypes, enabling effective end-to-end training in decentralized settings.
@@
  • We conduct extensive experiments on real-world image semantic segmentation datasets, demonstrating that FUSS combined with FedCC consistently outperforms local DNN training and extensions of traditional FL aggregation strategies in fully unsupervised i.i.d and non-i.i.d federated scenarios, while preserving scalability and privacy.
2 Unsupervised image Semantic Segmentation
USS aims to segment images into semantically coherent regions without requiring human labels. Recent methods follow a common two-stage pipeline: a) extract dense pixel-level feature embeddings using pretrained self-supervised models, and b) optimize over these representations using unsupervised objectives to produce segmentation masks.
@@ 2 Unsupervised image Semantic Segmentation
Feature extraction. Let $E$ denote a pretrained DNN encoder that maps an image $\mathbf{X}_{i}$ to a spatial feature tensor $\mathbf{H}_{i}=E(\mathbf{X}_{i})$, with $\mathbf{H}_{i}\in\mathbb{R}^{H\times W\times D}$. Each location $(h,w)$ in $\mathbf{H}_{i}$ corresponds to a $D$-dimensional embedding capturing local semantic context.
To enable clustering, the encoder must produce similar features for pixels depicting the same semantic object, while allowing different objects to be distinguishable. We define a general similarity function:
@@
where $\psi:\mathbb{R}^{D}\times\mathbb{R}^{D}\rightarrow\mathbb{R}$ is a base similarity metric (e.g., cosine similarity), and $\phi:\mathbb{R}\rightarrow\mathbb{R}$ is an optional normalization function. Higher values of $s$ indicate stronger semantic affinity between pixel representations, supporting clustering-based label discovery.
To guide the encoder $E$ towards learning semantically coherent pixel-wise representations, contrastive objectives are typically employed (Caron et al., 2021). The resulting loss functions promote intra-object compactness by increasing the similarity among perceptually related pixels, while simultaneously encouraging inter-object separability by reducing similarity between unrelated regions.
@@
Pixel feature clustering. Following feature extraction, the second stage of the USS pipeline involves grouping pixel-level feature vectors into $|\mathcal{C}|$ distinct clusters, each corresponding to a semantic region. We define a matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{C}|\times D}$, where each row $\mathbf{m}_{c}$ is a cluster centroid representing a semantic prototype for the $c$-th region or object class. To associate pixel embeddings with cluster prototypes, we first reshape the encoder output $\mathbf{H}_{i}\in\mathbb{R}^{H\times W\times D}$ into a matrix $\tilde{\mathbf{H}}_{i}\in\mathbb{R}^{(HW)\times D}$, where each row corresponds to the feature vector of a spatial location. We then compute the similarity matrix via inner product:
@@
Instead, the centroid matrix $\mathbf{M}$ is initialized randomly and incrementally optimized with each mini-batch of extracted features (MacQueen, 1967). The goal is for pixel representations assigned to the same centroid to be similar, while different centroids remain well-separated. This is typically achieved by minimizing intra-cluster variance and penalizing inter-cluster similarity:
@@ 3 Federated Unsupervised Semantic Segmentation (FUSS)
FUSS extends the standard Federated Learning (FL) paradigm (McMahan et al., 2017; Li et al., 2020, 2021) to the unsupervised semantic segmentation setting, where clients must collaboratively train segmentation models without any labeled data or data sharing. Let $\mathcal{K}$ denote the set of participating clients, each indexed by $k$. Each client holds a private unlabeled dataset $\mathcal{D}_{k}=\{\mathbf{X}_{k_{i}}\in\mathbb{R}^{H\times W\times 3}\mid i=1,\dots,N_{k}\},$ where $N_{k}$ is the number of local images. The global dataset is $\mathcal{D}=\bigcup_{k\in\mathcal{K}}\mathcal{D}_{k}$, with total size $N=\sum_{k\in\mathcal{K}}N_{k}$.
@@
Unlike traditional FL setups, FUSS must ensure consistency not only in feature extraction but also in semantic clustering. Aligning the learned prototypes $\mathbf{M}_{k}$ across clients is nontrivial, as potential local data heterogeneity can induce variation in learned feature spaces and disrupt semantic alignment. To mitigate the adverse effects of data heterogeneity and class imbalance during local representation learning, FUSS leverages recent advances in vision foundation models (VFMs) such as DINO (Caron et al., 2021), which produce semantically meaningful pixel embeddings without supervision, that coarsely align semantically similar regions even across unseen classes. Due to space constraints, a visual representation of the FUSS framework is deferred to Appendix A.

### v2

@@ Abstract
This work explores the application of Federated Learning (FL) to Unsupervised Semantic image Segmentation (USS). Recent USS methods extract pixel-level features using frozen visual foundation models and refine them through self-supervised objectives that encourage semantic grouping. These features are then grouped to semantic clusters to produce segmentation masks. Extending these ideas to federated settings requires feature representation and cluster centroid alignment across distributed clients, an inherently difficult task under heterogeneous data distributions in the absence of supervision. To address this, we propose FUSS (Federated Unsupervised image Semantic Segmentation) which is, to our knowledge, the first framework to enable fully decentralized, label-free semantic segmentation training. FUSS introduces novel federation strategies that promote global consistency in feature and prototype space, jointly optimizing local segmentation heads and shared semantic centroids. Experiments on both benchmark and real-world datasets, including binary and multi-class segmentation tasks, show that FUSS consistently outperforms local-only client trainings as well as extensions of classical FL algorithms under varying client data distributions. To fully support reproducibility, the source code, data partitioning scripts, and implementation details are publicly available at: https://github.com/evanchar/FUSS
keywords:
Federated Learning , Image Segmentation , Unsupervised Learning , Prototype Alignment , FedCC
journal: Neurocomputing
@@ 1 Introduction
Semantic image segmentation is a fundamental computer vision task, with the objective of assigning image pixels to pre-defined class labels. Given its central role in scene understanding, semantic segmentation serves as a core component of visual perception systems deployed in safety-critical applications like autonomous driving [36], industrial inspection [30], or medical diagnostics [5, 31]. Unsupervised Semantic Segmentation (USS) has emerged as a family of methods that do not require pixel-wise annotations, which are notoriously expensive. While the absence of manual annotations makes it infeasible to directly assign semantic labels to image regions, USS methods aim to discover coherent visual groupings whose emergent categories exhibit strong alignment with human-perceived semantic concepts. Recent USS approaches employ pretrained Convolutional Neural Networks (CNNs) [15] or Vision Foundation Models (VFMs) [4, 8] to extract pixel-wise feature embeddings. These representations are fed into learnable segmentation heads that are fine-tuned through self-supervised learning objectives such as contrastive losses [11, 13], energy minimization [14], or eigenvector alignment [18]. Unsupervised segmentation is ultimately achieved by simply clustering the learned representations.
In safety-critical domains, strict privacy constraints on data storage and sharing introduce significant challenges in effectively training robust and generalizable DNN-based image segmentation models. For instance, hospitals routinely collect large volumes of unlabeled radiology or pathology scans but are unable to share them across institutions due to privacy regulations such as HIPAA [17] or GDPR [38]. This restriction can lead to biased or under-trained segmentation models. Similarly, industrial facilities may require semantic segmentation models for pipeline and machinery visual inspection [30], yet they face both annotation scarcity and limitations on transmitting proprietary data, which may lead to undetected defects and increased operational risk. In the case of autonomous vehicles, edge devices must adapt segmentation models to local video streams without access to annotations [33] and under strict data-sharing constraints, where failure to do so can induce bias or silent model degradation, with potentially severe safety consequences.
@@
To this end, Federated Learning (FL) [27, 2, 10, 9, 39] has emerged as a promising paradigm that enables collaborative model training with decentralized data. Perhaps the most recognizable FL method is FedAvg [27], a method that simply aggregates the weights between federation clients. The simplicity and elegance of FedAvg render it applicable to almost every supervised [34] or unsupervised training [25, 23] scenario, making FedAvg the standard benchmark in most FL applications.
In semantic image segmentation, federated learning (FL) has thus far been investigated primarily under supervised training regimes [28, 9], or within domain adaptation frameworks that rely on large annotated datasets for source-domain pretraining [43, 35]. A key challenge in this context arises from two distinct forms of data heterogeneity. First, domain shift (feature distribution skew) that occurs as a natural characteristic of real-world deployments. For example, distinct visual sensors may capture different backgrounds or lighting conditions. Furthermore, the same semantic object may exhibit heavy appearance variation in different locations (e.g., a pristine pipeline in a modern refinery versus a damaged pipeline in an unsupported industrial site). Second, label distribution skew (class imbalance) that forces clients to discover semantic prototypes from disjoint class subsets. In unsupervised settings with a fixed “class discovery” budget, this form of heterogeneity can be particularly challenging, as the absence of specific classes, may force local models to assign “empty” prototype slots to random features or noise, disrupting alignment across federated clients.
Supervised objectives, as demonstrated in [28], can partially mitigate this issue by aligning representations of similar objects toward consistent target values across clients. However, these mechanisms are inherently inapplicable in the Federated Unsupervised Semantic Segmentation (FUSS) setting, where no supervised signal is available at any stage of the segmentation pipeline. Consequently, learning coherent and transferable semantic representations under these constraints remains an open and largely unexplored research problem.
@@
To bridge this gap, we propose Federated Unsupervised image Semantic Segmentation (FUSS), a novel framework for collaborative unsupervised training of image segmentation DNN models in decentralized settings. FUSS addresses both data annotation scarcity and privacy constraints by enabling clients to jointly learn DNN segmentation heads and class prototypes using unsupervised objectives and communication-efficient parameter exchanges.
Across these domains, a common challenge emerges: When an entity requires an AI system for high-risk, safety-critical tasks but lacks adequate, unbiased, or annotated data (due to privacy or time constraints), centralized training cannot be trusted to produce reliable results. In such scenarios, where centralized annotation or data sharing is infeasible, federated unsupervised training is not just practical, it may be the only viable solution.
@@
  • 1.
    We formally define the FUSS framework, which combines unsupervised representation learning with semantic prototype alignment across clients. FUSS enables efficient and stable federated image segmentation training without any reliance on annotated image data, a characteristic that allows seamless applicability in heavily privacy-constrained data settings.
  • 2.
@@
    We develop a novel aggregation strategy, namely Federated Centroid Clustering (FedCC), tailored to federated unsupervised segmentation. FedCC supports joint optimization of client-specific segmentation heads and global alignment of semantic class prototypes, enabling effective end-to-end training in decentralized settings, even with highly heterogeneous clients.
  • 3.
    We establish the first comprehensive benchmarks for FUSS on the Cityscapes and CocoStuff datasets and validate our methods on a real-world industrial pipeline segmentation dataset.
@@
Through extensive experimentation, we demonstrate that USS, combined with FedCC, consistently outperforms local DNN training and extensions of traditional FL aggregation strategies in fully unsupervised federated scenarios, while preserving scalability and data privacy.
The remainder of this paper is structured as follows. Section 2 provides an overview of related work in unsupervised semantic image segmentation and federated learning. Section 3 introduces USS unified notation. Section 4 presents the proposed FUSS framework, detailing our methodology, architectural choices and aggregation strategies. Section 5 describes the experimental setup, including dataset partitions and federated configurations, and reports quantitative results along with key observations. Finally, Section 6 concludes the paper and outlines directions for future research.
2 Related Work
2.1 Unsupervised Semantic Segmentation
Early Unsupervised Semantic Segmentation (USS) approaches were not DNN-based [29]. Instead, they relied on low-level hand-crafted image features, such as color histograms, texture, and intensity gradients, to cluster pixels into coherent groups. While effective for simpler scenes, these methods often struggled with the complex semantic variability found in unrestricted real-world environments.
To mitigate this, earlier deep learning approaches introduced end-to-end unsupervised objectives to train Convolutional Neural Networks (CNNs). For instance, W-Net [42] concatenated two U-Net architectures into an autoencoder, jointly optimizing a soft normalized cut loss for segmentation and a reconstruction loss for data fidelity. DeepCluster [3] proposed an iterative framework where features from a CNN were clustered via k-means to generate pseudo-labels, which then supervised the network’s weight updates, effectively bootstrapping semantic representations.
More recently, PiCIE [6] advanced this direction by employing invariance to photometric transformations and equivariance to geometric shifts as inductive biases, enabling the clustering of pixels into semantically coherent groups even in non-object-centric datasets. However, despite these advancements, such methods often required extensive training from scratch and struggled to capture the rich, open-world semantic granularity offered by modern Vision Foundation Models.
With the emergence of self-supervised representation learning [12], USS saw significant progress. STEGO (Self-supervised Transformer with Energy-based Graph Optimization) [14] was a key breakthrough, which used self-supervised Vision Transformers [4] pre-trained on contrastive objectives to generate semantically structured feature maps, where pixels belonging to similar objects or regions within an image are mapped to proximate locations in the embedding space.
Building on this, EAGLE (Eigen Aggregation Learning for Object-Centric Unsupervised Semantic Segmentation) [18] introduced a spectral image segmentation approach that utilizes structural cues derived from the image color affinity graph to produce more object-centric and boundary-aware image segmentation. PriMaPs [13] introduced a novel paradigm that decomposes images into semantically meaningful “principal mask proposals” based on feature representations. By fitting class prototypes to these proposals via a stochastic expectation-maximization algorithm (PriMaPs-EM), the method effectively boosts unsupervised segmentation performance.
While these approaches offer refinements in specific scenarios, they often introduce additional computational overhead for incremental gains. STEGO remains a cornerstone of the field, providing highly competitive performance with significantly lower complexity. This balance of performance and efficiency motivates our adoption of STEGO as the core segmentation engine for our federated framework.
Although distinct in their objectives, the adjacent fields of Unsupervised Salient Object Detection (USOD) and Instance Segmentation have recently proposed mechanisms highly transferable to the USS context. In the domain of USOD, Liu et al. [24] introduced a Belief Capsule Network (BCNet) that captures object-part hierarchies, enabling the generation of refined pseudo-masks with enhanced object integrity. Moreover, for instance segmentation in data-scarce regimes, Zhang et al. [44] leveraged Vision-Language Models to generate semantically rich pseudo-masks, which subsequently serve as prompts for unsupervised pre-training of Query-based End-to-End Image Segmentation models (QEIS). These innovations introduce strong inductive biases through hierarchical grouping and language-driven guidance, that inspire directions for overcoming the granularity and initialization challenges inherent in unsupervised scene understanding.
Despite architectural and optimization differences, SOTA USS methods share a common foundation that also motivates our approach: the use of strong self-supervised DNN backbones for feature extraction, followed by unsupervised optimization techniques to uncover semantically meaningful groupings in the data.
2.2 Federated Learning
Federated Learning (FL) [27] has emerged as the standard for collaborative model training under privacy constraints, enabling distributed optimization without centralizing raw data. In the specific domain of semantic segmentation, existing FL frameworks have primarily focused on supervised settings, where clients are assumed to possess pixel-perfect ground-truth labels, but nontheless appear to be statistically heterogeneous, and attribute inherent in distributed environments.
To address this challenge, methods like FedDrive [9] introduced effective benchmarking strategies for domain generalization, utilizing style transfer to bridge the domain shift across autonomous driving datasets. Similarly, FedSeg [28] tackled high label distribution skew by correcting client local optimization with a modified cross-entropy loss and employing pixel-level contrastive learning to align local embeddings with a global semantic space.
While effective, these methods remain fundamentally constrained by their reliance on extensive annotated data, which is often unavailable or prohibitively expensive to acquire in real-world edge scenarios.
Parallel efforts have explored Unsupervised Federated Learning (UFL) to eliminate the dependency on labels. Approaches such as Orchestra [25] and FedU [45] have demonstrated that global clustering mechanisms and divergence-aware aggregation can yield robust representations even from non-IID data. Liao et al. [23] further advanced this by addressing representation collapse through a flexible uniform regularizer that preserves uniformity in the embedding space. However, these works predominantly target image-level classification tasks. They lack the dense, spatial granularity required to enforce semantic consistency at the pixel level, leaving the challenge of Federated Unsupervised Semantic Segmentation largely unaddressed.
To the best of our knowledge, FUSS is the first framework specifically designed to bridge this gap. Unlike previous methods that optimize for global image categories, FUSS introduces a novel paradigm for jointly aligning pixel-level feature representations and global cluster centroids in a fully decentralized, label-free manner. By extending the principles of self-supervised grouping to the federated setting, our approach enables the discovery of coherent semantic regions across heterogeneous clients without sharing raw data or annotations.
3 Preliminaries
@@
Feature extraction Let $E$ denote a pretrained DNN encoder that maps an image $\mathbf{X}_{i}$ to a dense feature representation $\mathbf{H}_{i}=E(\mathbf{X}_{i})$, with $\mathbf{H}_{i}\in\mathbb{R}^{H\times W\times D}$. Each location $(h,w)$ in $\mathbf{H}_{i}$ corresponds to a $D$-dimensional embedding capturing local semantic context.
To enable effective clustering, the encoder must generate feature representations in which pixels belonging to the same semantic object exhibit high similarity, while pixels from different objects remain clearly separable. We formalize this by defining a general similarity function:
@@
where $\psi:\mathbb{R}^{D}\times\mathbb{R}^{D}\rightarrow\mathbb{R}$ denotes a base similarity metric, such as cosine similarity, and $\phi:\mathbb{R}\rightarrow\mathbb{R}$ represents an optional normalization or scaling function. Tuple $(m,n)$ represents the spatial coordinates (row and column indices) of a second pixel in the feature map $\mathbf{H}$, distinct from the reference pixel at $(h,w)$. This general formulation permits the adoption of various similarity functions within the USS framework. Larger values of $s$ correspond to stronger semantic affinity between pixel-level representations, whereas values closer to the lower bound of the function indicate greater dissimilarity between feature representations. Such a formulation allows for effective clustering-based discovery of coherent object groupings.
To guide the encoder $E$ towards learning semantically coherent pixel-wise representations, contrastive objectives are typically employed [4]. The resulting loss functions promote intra-object compactness by increasing the similarity among perceptually related pixels, while simultaneously encouraging inter-object separability by reducing similarity between unrelated regions.
@@
Pixel feature clustering Following feature extraction, the second stage of the USS pipeline involves grouping pixel-level feature vectors into $|\mathcal{C}|$ distinct clusters, each corresponding to a semantic region. We define a matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{C}|\times D}$, where each row $\mathbf{m}_{c}$ is a cluster centroid representing a semantic prototype for the $c$-th region or object class. To associate pixel embeddings with cluster prototypes, we first reshape the encoder output $\mathbf{H}_{i}\in\mathbb{R}^{H\times W\times D}$ into a matrix $\tilde{\mathbf{H}}_{i}\in\mathbb{R}^{(HW)\times D}$, where each row corresponds to the feature vector of a spatial location. We then compute the similarity matrix via inner product:
@@
Instead, the centroid matrix $\mathbf{M}$ is initialized randomly and incrementally optimized with each mini-batch of extracted features [26]. The goal is for pixel representations assigned to the same centroid to be similar, while different centroids remain well-separated. This is typically achieved by minimizing intra-cluster variance and penalizing inter-cluster similarity:
@@ 4 Federated Unsupervised Semantic Segmentation (FUSS)
FUSS extends the standard Federated Learning (FL) paradigm. [27, 22, 21] to the unsupervised semantic segmentation setting, where clients must collaboratively train segmentation models without any labeled data or data sharing.
Figure 1: Overview of the proposed FUSS framework. Each client performs local unsupervised training using self-supervised objectives and periodically transmits its learned segmentation head and semantic prototypes to a central server. The server aggregates these parameters and broadcasts the updated global model back to all clients. While traditional federated strategies may result in disorganized or overlapping class prototypes due to local heterogeneity, our proposed aggregation methods explicitly promote semantic alignment and prototype separability across the federation.
Let $\mathcal{K}$ denote the set of participating clients, each indexed by $k$. Each client holds a private unlabeled dataset $\mathcal{D}_{k}=\{\mathbf{X}_{k_{i}}\in\mathbb{R}^{H\times W\times 3}\mid i=1,\dots,N_{k}\},$ where $N_{k}$ is the number of local images. The global dataset is $\mathcal{D}=\bigcup_{k\in\mathcal{K}}\mathcal{D}_{k}$, with total size $N=\sum_{k\in\mathcal{K}}N_{k}$.
@@
Unlike traditional FL setups, FUSS must ensure consistency not only in feature extraction but also in semantic clustering. Aligning the learned prototypes $\mathbf{M}_{k}$ across clients is nontrivial, as potential local data heterogeneity can induce variation in learned feature spaces and disrupt semantic alignment. To mitigate the adverse effects of data heterogeneity and class imbalance during local representation learning, FUSS leverages recent advances in vision foundation models (VFMs) such as DINO [4], which produce semantically meaningful pixel embeddings without supervision that coarsely align semantically similar regions even across unseen classes.